# Single color virtual H&E staining with In-and-Out Net

### v1

@@ 3 Results and discussion
3.1 Model comparison
In Fig. 3, we can observe example outputs from the different models tested.
@@ 3.1 Model comparison
In Table. 1, we utilized MSE and PSNR to assess the model outputs across all test group images. These metrics provide valuable insights into the statistical similarity between the output images and their corresponding ground truth images. It is worth noting that In-and-Out Net emerged with the lowest MSE, suggesting a high degree of accuracy in its predictions. Conversely, VSGD-Net showcased superior performance in PSNR, indicating its ability to preserve signal fidelity amidst noise and distortion. However, in the context of virtual staining, we focus more on capturing structural details accurately. Structural similarity becomes more appropriate in determining the efficacy of the models. To this end, we employed SSIM, FSIM, and MS-SSIM to comprehensively evaluate the models’ ability to preserve structural integrity. Notably, In-and-Out Net outperformed all other models in the assessment of structural similarity. This suggests its exceptional capability in capturing the intricate details crucial for virtual staining applications.
@@ 3.2 Ablation study
In Table. 3, we systematically removed each key component of the In-and-Out Net to evaluate their individual contributions. Initially, we conducted an ablation study on the In-and-Out Training, which involved training the inner and outer loops alternately. For this ablation, we trained all components simultaneously while setting $\alpha=0.5$ in Eq. 8, denoted as Ablation 1. Subsequently, we performed ablations on $D_{out}$ and $D_{H}\&D_{E}$ separately, referred to as Ablation 2 and Ablation 3, respectively. Finally, we removed the branches at the beginning of the model architecture, where the RCM images were trained into two branches: H and E channel, labeled as Ablation 4. The last row in Table. 3 represents the performance of the original In-and-Out Net without any ablations.
@@
Table. 3 shows that the original model outperforms all ablation cases. Notably, for Ablation 3, the VOL value is significantly small, leading us to disregard its MSE and SSIM metrics due to the observed blurriness. This is reasonable because $D_{H}\&D_{E}$ were applied on H and E channel generated images during inner loop training, which focuses on learning detailed structure information. So, after ablating $D_{H}\&D_{E}$, the output images become blurry.
Fig. 5 illustrates examples of outputs from the ablation study. In Ablation 1, where In-and-Out training was omitted, the images exhibit slight blurriness, and the learned structures appear less defined. Ablation 2, which involved removing $D_{out}$, resulted in faded colors and a little ambiguous structure boundaries. In Ablation 3, where $D_{H}\&D_{E}$ were removed, the images are noticeably blurry with minimal structure details. As for Ablation 4, which mirrors the setup of pix2pixGAN, we have not included examples from this ablation in Fig. 5.
@@ 3.3 Training approaches
Fig.6 illustrates the impact of different training approaches on $eval\_loss$, as defined in Eq.10. Our training methodology involves alternating between $n$ epochs of inner loop training followed by $n$ epochs of outer loop training, and repeating this cycle until a total of $400$ epochs have been completed. In Fig. 6, the values of $n$ are indicated in the legend. Additionally, we include the results of Ablation 1 in the figure for comparison, as it was trained without the In-and-Out training approach.
@@ 3.4 Discussion
In-and-Out Net effectively transforms RCM images into H&E-stained images with remarkable color fidelity and accurate structural representation. Unlike other models, it is purposefully designed for virtual staining tasks. In this study, we produced H&E ground truth images comprising two distinct fluorescence channels, enabling the separate training of both channels. However, even with real H&E-stained images as ground truth, we can utilize color decomposition to isolate the H and E channel images for employment with In-and-Out Net. The good performance of In-and-Out Net comes from its utilization of the prior knowledge that H&E-stained images consist of two fixed colors. Moreover, it employs a dual-stage training approach, comprising coarse training (inner loop) and fine training (outer loop), to effectively converge towards an optimal local minimum for the loss function.
A limitation of our study is the inadequate quality of RCM images. As shown in Table. 1, even the highest SSIM attained is merely 0.5081. This hampers our ability to morphologically compare specific structures accurately, primarily due to the poor quality of RCM images utilized. In most studies, investigators would employ commercial RCM systems such as the VivaScope 1500 System (Caliber I.D., Rochester, NY) for imaging, which offer superior Signal-to-Noise Ratio (SNR), enhanced contrast, superior resolution, and more detailed structural information compared to our dataset. One potential explanation could be the disparity in laser wavelengths, as we used a 488 nm laser for RCM imaging, while the Caliber system utilizes an 830 nm laser. Additionally, the system we employed is not specifically designed for RCM, leading to some optical limitations for reflectance mode. Owing to these system limitations, acquiring high-quality RCM data is challenging, which results in most images captured from dermal/epidermal junctions. So, our analysis lacks thorough examination within different tissue categories. Nonetheless, these limitations can be addressed with improved RCM inputs, and they do not undermine the efficacy of our In-and-Out Net methodology. While we’ve shown that creating ground truth with two fluorescence images is feasible, it’s essential for clinical experts to validate this approach in the future. This validation will help determine if there are significant differences in information between the pseudo ground truth and actual H&E staining images.
@@ 4 Conclusion
In this study, we introduce a novel virtual staining network called In-and-Out Net. We utilize this network to virtually stain RCM images to H&E-stained images and compare its performance with other networks. Additionally, we conduct an ablation study to validate the functionality of each component within our network. Furthermore, we compare various training approaches. Through the utilization of In-and-Out Net, we achieve successful staining of RCM images to H&E with excellent color and structural matching, outperforming other models. Through the ablation study, we confirm the importance of In-and-Out training and validate the functions of the inner loop and outer loop by ablating three discriminators. Moreover, we investigate the effects of different training approaches. Despite the limitations imposed by the quality of RCM images, we are the first to virtually stain aluminum-processed RCM images to H&E images. We anticipate that In-and-Out Net can be widely applied to various virtual staining tasks.

### v2

@@ 3 Results and discussion
3.1 Model results comparison
In Figure 3, we can observe example outputs from the different models tested.
@@ 3.1 Model results comparison
In Table 2, we utilized MSE and PSNR to assess the model outputs across all test group images. These metrics provide valuable insights into the statistical similarity between the output images and their corresponding ground truth images. It is worth noting that In-and-Out Net emerged with the lowest MSE, suggesting a high degree of accuracy in its predictions. Conversely, VSGD-Net showcased superior performance in PSNR, indicating its ability to preserve signal fidelity amidst noise and distortion. However, in the context of digital staining, we focus more on capturing structural details accurately. Structural similarity becomes more appropriate in determining the efficacy of the models. To this end, we employed SSIM, FSIM, and MS-SSIM to comprehensively evaluate the models’ ability to preserve structural integrity. Notably, In-and-Out Net outperformed all other models in the assessment of structural similarity. This suggests its exceptional capability in capturing the intricate details crucial for digital staining applications.
@@
As for traditional machine learning methods, we would like to clarify that no comparison with such methods was included because, as far as we know, traditional machine learning models are not well-suited for solving the digital staining problem. These models typically excel in classification and regression tasks but lack the ability to perform pixel-wise image-to-image translations that are necessary for generating digital stained images. Furthermore, traditional machine learning methods have not been shown in the literature to successfully address the complex transformations required for digital staining. As such, no prior work has applied traditional machine learning to this task, making deep learning models, particularly Generative Adversarial Networks (GANs), the most effective and appropriate choice for this application.
@@ 3.2 Ablation study
In Table 4, we systematically removed each key component of the In-and-Out Net to evaluate their individual contributions. Initially, we conducted an ablation study on the In-and-Out Training, which involved training the inner and outer loops alternately. For this ablation, we trained all components simultaneously while setting $\alpha=0.5$ in Equation 8, denoted as Ablation 1. Subsequently, we performed ablations on $D_{out}$ and $D_{H}\&D_{E}$ separately, referred to as Ablation 2 and Ablation 3, respectively. Finally, we removed the branches at the beginning of the model architecture, where the RCM images were trained into two branches: H and E channel, labeled as Ablation 4. The last row in Table 4 represents the performance of the original In-and-Out Net without any ablations.
Table 4 shows that the original model outperforms all ablation cases. Notably, for Ablation 3, the VOL value is significantly small, leading us to disregard its MSE and SSIM metrics due to the observed blurriness. This is reasonable because $D_{H}\&D_{E}$ were applied on H and E channel generated images during inner loop training, which focuses on learning detailed structure information. So, after ablating $D_{H}\&D_{E}$, the output images become blurry.
Figure 5 illustrates examples of outputs from the ablation study. In Ablation 1, where In-and-Out training was omitted, the images exhibit slight blurriness, and the learned structures appear less defined. Ablation 2, which involved removing $D_{out}$, resulted in faded colors and a little ambiguous structure boundaries. In Ablation 3, where $D_{H}\&D_{E}$ were removed, the images are noticeably blurry with minimal structure details. As for Ablation 4, which mirrors the setup of pix2pixGAN, we have not included examples from this ablation in Figure 5.
3.3 Model complexity comparison
Table 5 presents a comparison of model complexity across various methods, including FLOPs (Floating Point Operations), number of parameters, inference time, and training time. The In-and-Out Net has a higher number of parameters and FLOPs comapred to other models except pix2pixHD, which is due to its dual-branch structure, but still maintains a feasible inference time, making it applicable for clinical use. Notably, while VSGD-Net and pix2pixHD have a similar number of parameters to our model, we achieve a shorter inference time and significantly reduced training time.
A larger number of parameters generally provides greater capacity for learning from data, which is why we additionally provide a complexity comparison across the four ablation studies in Section 3.2. According to the table, Ablation 1, 2, and 3 have a similar number of parameters as the In-and-Out Net, yet Table 4 clearly shows that In-and-Out Net achieves significantly better performance. The ablation studies highlight the effect of different components on model complexity and performance, demonstrating how each part contributes to the overall efficiency and accuracy of the model. This comparison illustrates that while In-and-Out Net has increased complexity, its performance gains are not merely due to the larger parameter count but rather the innovative architecture and training strategy employed. In addition, the inference time for a single image is sufficiently low for clinical use, meaning that despite the increased complexity of our model, it is not prohibitively high to interfere with the usability of it.
Table 5: Model complexity comparison across different methods. FLOP: Floating Point Operation. The table reports the FLOPs (in GFLOPs), number of parameters (in millions), inference time (in milliseconds), and training time (in hours) for each model. The In-and-Out Net shows increased complexity due to the dual-branch structure, but maintains an acceptable inference time for clinical applications. Ablation studies are included to demonstrate the contribution of different model components to the overall complexity and performance.
Method FLOPs (Inference) (GFLOPs) Number of Parameters (Million) Inference Time (ms) Training Time (hour)
pix2pixGAN 13.8 55.3 254 8.01
pix2pixHD 54.4 182 547 $\geq$24
VSGD-Net 33.7 91.4 1551 $\geq$48
RestainNet loss 13.8 55.3 252 8.11
Ablation 1 27.7 111 450 16.0
Ablation 2 27.7 111 450 15.5
Ablation 3 27.7 110 450 14.7
Abaltion 4 13.8 55.3 254 8.01
In-and-Out Net 27.7 111 450 16.2
@@ 3.4 Training approaches
Figure 6 illustrates the impact of different training approaches on $eval\_loss$, as defined in Equation 10. Our training methodology involves alternating between $n$ epochs of inner loop training followed by $n$ epochs of outer loop training, and repeating this cycle until a total of $400$ epochs have been completed. In Figure 6, the values of $n$ are indicated in the legend. Additionally, we include the results of Ablation 1 in the figure for comparison, as it was trained without the In-and-Out training approach.
@@ 3.5 Discussion
In-and-Out Net effectively transforms RCM images into H&E-stained images with remarkable color fidelity and accurate structural representation. Unlike other models, it is purposefully designed for digital staining tasks. In this study, we produced H&E ground truth images comprising two distinct fluorescence channels, enabling the separate training of both channels. However, even with real H&E-stained images as ground truth, we can utilize color decomposition to isolate the H and E channel images for employment with In-and-Out Net. The good performance of In-and-Out Net comes from its utilization of the prior knowledge that H&E-stained images consist of two fixed colors. Moreover, it employs a dual-stage training approach, comprising coarse training (inner loop) and fine training (outer loop), to effectively converge towards an optimal local minimum for the loss function.
A limitation of our study is the inadequate quality of RCM images. As shown in Table 2, even the highest SSIM attained is merely 0.5081. This hampers our ability to morphologically compare specific structures accurately, primarily due to the poor quality of RCM images utilized. In most studies, investigators would employ commercial RCM systems such as the VivaScope 1500 System (Caliber I.D., Rochester, NY) for imaging, which offer superior Signal-to-Noise Ratio (SNR), enhanced contrast, superior resolution, and more detailed structural information compared to our dataset. One potential explanation could be the disparity in laser wavelengths, as we used a 488 nm laser for RCM imaging, while the Caliber system utilizes an 830 nm laser. Additionally, the system we employed is not specifically designed for RCM, leading to some optical limitations for reflectance mode. Owing to these system limitations, acquiring high-quality RCM data is challenging, which results in most images captured from dermal/epidermal junctions. So, our analysis lacks thorough examination within different tissue categories. Nonetheless, these limitations can be addressed with improved RCM inputs, and they do not undermine the efficacy of our In-and-Out Net methodology. While we’ve shown that creating ground truth with two fluorescence images is feasible, it’s essential for clinical experts to validate this approach in the future. This validation will help determine if there are significant differences in information between the pseudo ground truth and actual H&E staining images.
@@ 4 Conclusion
In this study, we introduce a novel digital staining network called In-and-Out Net. We utilize this network to digitally stain RCM images to H&E-stained images and compare its performance with other networks. Additionally, we conduct an ablation study to validate the functionality of each component within our network. Furthermore, we compare various training approaches. Through the utilization of In-and-Out Net, we achieve successful staining of RCM images to H&E with excellent color and structural matching, outperforming other models. Through the ablation study, we confirm the importance of In-and-Out training and validate the functions of the inner loop and outer loop by ablating three discriminators. Moreover, we investigate the effects of different training approaches. Despite the limitations imposed by the quality of RCM images, we are the first to digitally stain aluminum-processed RCM images to H&E images. We anticipate that In-and-Out Net can be widely applied to various digital staining tasks.